what are the x and y coordinates of y=-3x+1?
thanks

Answer by Abbey(339)

x and y coordinates are determined by the values you plug into the equation. The y is the dependant variable and the x is the independant variable. In other words, the value for y is based on what you select for x.
If you select x=0, then y will be 1, you would say the x and y coordinates are (0,1)
y=-3x+1
y=-3(0)+1
y=0+1
y=1
It is completely up to you which number you want to plug in for the x value.
If you chose x=1, then y would be -2 and your coordinate would be (1,-2)
because
y=-3x+1
y=-3(1)+1
y=-3+1
y=-2